Financial Accounting questions multiple choice about the chapter ( Cash & Investments )

1)
On September 30, 20x6, Riches Inc. made three investments: 1221 shares of Wendle Inc., 4175 shares of Sam International, and 823 shares of Hinter Inc. The shares of Wendle Inc. and Hinter Inc. are classified as FVTPL and the shares of Sam International are classified as FVTOCI.
The market price per share is as follows:

Wendle Inc.
Sam International
Hinter Inc.
Sept. 30, x6
$12.85
$5.73
$7.90
Dec. 31, x6
$16.03
$8.20
$4.00

What amount would be reported as total comprehensive income for the year 20x6, assuming income before gains/losses on investments is $141458?

Select one:
a. $10312
b. $142131
c. $10985
d. $152443

2)The bank reconciliation for Julia’s company showed an adjusted cash balance of $6527 as the correct May 31st 20x4 balance. During May 20x4, the bank statement was received and included the following:

Bank service charges
$184
Interest on bank loan
$213

In addition, during May 20x4 $2546 of outstanding cheques had not yet cleared the bank. What is the unadjusted (i.e. before receiving the bank statement) cash balance at May 31st, 20x4?

Select one:
a. $6924
b. $6130
c. $9470
d. $3584

3)Spector Ltd. had a cash balance of $17059 on the books at the end of December 20x10. The following items were noted by the controller at Spector after comparing the bank statement and the company’s cash account:

A customer’s returned cheque (NSF) was included with the bank statement. The cheque was for $3819.
Monthly service charges of $222 were included in the bank statement.
Outstanding cheques at the end of December totaled $2353.
A December 31, 20x10 deposit of $7957 had not been received by the bank in time to be included in the December bank statement.

Which of the following is the correct revised cash balance at December 31, 20x10?


Select one:
a. $18622
b. $13018
c. $10665
d. $28613

4)A company purchased shares costing $108199 during the year. These shares are classified as FVTOCI. At the end of the year, the company received $3562 in dividends from these shares. At year-end, the fair value of the shares is $167719. What is the net impact on the Statement of Comprehensive Income for the year?
Select one:
a. $3562 revenue in profit and loss and $0 in other comprehensive income
b. $3562 revenue in profit and loss and $59520 in other comprehensive income
c. $0 revenue in profit and loss and $63082 in other comprehensive income
d. $63082 revenue in profit and loss and $0 in other comprehensive income


5)The accountant for Pardy Inc. gathered the following information to prepare the December 31, 20x3 bank reconciliation:


1. Cheques amounting to $5974 are outstanding at December 31

2. The December 31, 20x3 deposit of $2564 had not been received by the bank in time to be included in the December bank statement

3. The bank charged the company $230 for monthly standard service charges and have not yet been recorded by the company

4. The balance in Pardy Inc.’s cash account according to its accounting records was $9546


What is the cash balance per bank at December 31, 20x3?

Select one:
a. $5906
b. $15290
c. $9316
d. $12726
Answer #1

Part 1

Answer is option D

D. $152443

Fair value adjustment = (1221*(16.03-12.85))+(4175*(8.20-5.73))+(823*(4.00-7.90)) = $10985

Total comprehensive income for the year 20x6 = 141458+10985 = $152443

Part 2

Answer is option A

A. $6924

Unadjusted cash balance at May 31st, 20x4= adjusted cash balance + bank service charges + interest on bank loan = 6527+184+213 = $6924

Part 3

Answer is option B

B. $13018

Adjusted cash balance = unadjusted cash balance – NSF cheque – bank service charges = 17059-3819-222 = $13018

Part 4

Answer is option B

B. $3562 revenue in profit and loss and $59520 in other comprehensive income

Other comprehensive income = (167719-108199) = 59520

Part 5

Answer is option C

C. $9316

Cash balance per bank at December 31, 20x3 = cash balance per bank – bank charges = 9546-230 = $9316